OpenStudy (anonymous):

Whats the value of tan 11π/6?

OpenStudy (anonymous):

11^(Pi/6) or 11 multiplied by pi/6??

OpenStudy (anonymous):

11^(pi/6)=3.510 or 11(Pi/6)=5.760

OpenStudy (anonymous):

tan 11pi divided by six on a \[\tan \frac{ 11\pi }{ 6 }\] unit sircle

OpenStudy (anonymous):

Hint 11pi/6 has a reference angle of pi/6 or 30 degrees. Is tangent negative of positive in the 4th quadrant?

Parth (parthkohli):

\[\rm \tan \left(11\pi \over 6 \right) = \tan\left( 11 {\pi \over 6}\right) = \tan \left(11\cdot 30 \right) = \tan \left(330^{\circ} \right) = {{\sin\left(330^{\circ}\right) } \over\cos(330^{\circ}) }\]

OpenStudy (anonymous):

\[\text{Tan}\left[\frac{11\pi }{6}\right]=-\frac{1}{\sqrt{3}} \]